OpenStudy (anonymous):

In a restaurant Patty's share of the expenses for lunch was $2 more than Karen's. Karen did not have enough money to pay her share, so Patty paid $8 of the expenses and Karen paid the remaining $4. How much did Karen owe Patty to settle the lunch expenses? 1. $1 2. $3 3. $5 4. $6 5. $9

OpenStudy (a_clan):

P - K = 2 .....(1) P + K = (8+4) =12 .......(2) On solving, P = $7 and K = $5 Since P paid $8 , so K owes $1 to P. Answer (1)

OpenStudy (anonymous):

thx